Mathematics 242 – Principles of Analysis Definitions Quiz 7 – April 4, 2014

Your Name:

B) (5) Let f be defined on an interval containing c. Define: f is differentiable at c.

A) B) (5) Let

f(x) = \{ \begin{matrix}1 − 3x^{2} & if x < 0 \\ 1 + 2x^{2} & if x ≥ 0. \end{matrix}

Is f differentiable at c = 0? Why or why not?